the length of a poster is 10cm longer than its width. if the area of the poster is 1200cm square, what is the width of the poster.
Answer by TimothyLamb(4379)

L = w + 10
Lw = 1200
---
Lw = 1200
(w + 10)w = 1200
ww + 10w - 1200 = 0
---
the above quadratic equation is in standard form, with a=1, b=10, and c=-1200
---
to solve the quadratic equation, plug this:
1 10 -1200
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
Answer 1:
the quadratic equation has two real roots:
root1 = 30
root2 = -40
---
the negative root doesn't make sense for a length, so use the positive root.
w = 30 cm
L = 40 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php